OpenStudy (anonymous):

An object is at the end of a spring at the bottom of the ramp.The spring is stretched 2.5 cm.What is the horizontal range of the object? k=1500 N/m,incline =30 degrees)

OpenStudy (priyar):

do u have a diagram that goes with this?

OpenStudy (priyar):

@iceman60 ?

OpenStudy (anonymous):

This is what I madee. http://imgur.com/kmIHk7e .No diagram is available

OpenStudy (michele_laino):

Hint: on the object are acting two forces: 1) the force due to gravity, whose magnitude is \(mg \sin \theta\) where \(\theta =30\) degrees 2) the eleastic force, whose magnitude is \(K \delta\), where \(\delta =2.5 \cdot 10^{-2}\) meters so, at equilibrium, we have: \[\Large mg\sin \theta = K\delta \] And the horizontal range is gi: \(\Large 2 \delta \cos \theta\)